# Conversational AI increases political knowledge as effectively as self-directed internet search

### v1

@@ Results
Figure 2: Conversational AI usage patterns and influence on belief in true versus false information. (A) Survey results: Self-reported use cases for AI chatbots among UK users. (B) RCT results: Change in agreement with true (purple) vs. false information (orange) from pre to post researching. Left panel shows the conversational AI condition; right panel shows the internet search control condition. Solid lines indicate researched topics, dotted lines denote non-researched topics. Error bars represent standard error of the mean. (C) RCT results, left: Bayesian GLM parameter estimates, error bars denote Highest Posterior Density Interval (HPDI). Gray shaded area depicts an apriori defined region of practical equivalence (ROPE), where effect sizes are considered to be negligible/practically 0; Right: GLM comparison using Widely Applicable Information Criterion (WAIC), as a measure of out-of-sample predictive accuracy of the GLMs. Full model = GLM1: GLM including parameters to quantify differences in change effects between conversational AI and internet search conditions, No ConvAI Term = GLM2: GLM not including parameters to quantify differences between different conversational AI models, Null model = GLM3: GLM not including parameters to quantify differences in change effects between conversational AI and internet search conditions or different conversational AI models
Randomised Controlled Trial (RCT)
@@
In a second study, independent of the survey sample, we recruited a separate sample of UK residents (N = 1,147 final sample) online via Prolific. Participants in the RCT researched true or false information related to issues of concern for UK voters (climate change, immigration, criminal justice, COVID-19 policy; see Supplementary Information for details on sourcing and political balancing of the material), either using conversational AI (GPT-4o, Claude-3.5, or Mistral) or internet search engines (Fig. 1). In separate studies, we randomly assigned participants to conduct research using a conversational AI model (GPT-4o) that was either instructed with a default prompt (baseline/control condition) or specifically prompt-engineered to be persuasive or sycophantic (treatment condition). After completing the study, participants were fully debriefed about the aims and hypotheses of the research. We assessed beliefs in true and false information on a 7-point Likert scale before and after researching all topics. Additionally, we measured secondary outcomes: trust, private political beliefs and extremism change.
@@
To test our research questions, we fitted and compared three Bayesian Generalized Linear Models (GLM1-3, Eq. 1, see Methods for details). There was no model evidence of differences between conversational AI and internet search conditions. Model fit metrics suggested no better fit of GLMs that included parameters for differences in change effects between conversational AI and internet search conditions (GLM1: Full model) in comparison to a GLM that did not include these terms (GLM3: Null model) (Fig. 2C, right panel). True information received on average approximately one Likert scale point higher agreement ratings than false information (Fig. 2B, purple vs orange lines), resulting in a non-zero difference parameter in GLM2 ($\beta_{\mathrm{TRUE}}=.67$, 95%-Highest Posterior Density Interval (HPDI) [.65; .69] (Fig. 2C right panel). Researching (vs not researching) political issues increased belief in true information and decreased belief in misinformation across time points (Fig. 2B, solid vs dotted lines; ($\beta_{\mathrm{POST}\times\mathrm{TRUE}\times\mathrm{RESEARCHED}}=.26$, 95%-HPDI [.19; .32], Fig. 2C). Importantly, we found that belief change was nearly identical for participants who researched using conversational AI or internet search engines (Fig. 2B, left vs right panel), reflecting in a close to zero parameter estimate ($\beta_{\mathrm{POST}\times\mathrm{TRUE}\times\mathrm{RESEARCHED}\times\mathrm{CONVAI}}=.02$, 95%-HPDI [–.09; .13], Fig. 2C). This null difference between conditions was further qualified by the fact that the HPDI fully encloses a region of practical equivalence, (ROPE; gray shaded zone), an apriori specified interval of effect sizes that are negligible. This pattern also held separately for each of the different model families tested (GPT, Claude, Mistral).
@@
We repeated the above analyses for trust, private political beliefs and extremism change, and found highly similar results as for beliefs in true and false information. Trust was measured by asking participants to state their agreement with statements on trust or distrust in institutions, experts, media and technology (see Supplementary Information for details). Private political beliefs were assessed by asking participants to state their private political beliefs for the 4 topics (see Supplementary Information for details). Extremism change was defined as sign flips in private political beliefs from before to after researching/not researching an issue – with referenced to the center point of the Likert scale (3.5, values below this value being negative, and values above being positive).
@@
For trust, we found that within GLM2, true statements on average produced higher ratings than false statements ($\beta_{TRUE}=.36$, 95%HPDI [.33; .39]). There was no change of trust ratings from before to after researching topics ($\beta_{POSTxTRUExRESEARCHED}=-.08$, 95%-HPDI [–.17; .009]). Importantly, we found that the patterns of trust and trust change were nearly identical for participants who researched using conversational AI or internet search engines ($\beta_{POSTxTRUExRESEARCHEDxCONVAI}=.026$, 95%-HPDI [–.08; .14]). For private political beliefs, true statements on average produced higher agreement ratings than false statements ($\beta_{TRUE}=.66$, 95%HPDI [.63; .70]). Additionally, across measurement time points, researching topics increased agreement with issues participants supported and increased disagreement with with issues participants did not support ($\beta_{POSTxTRUExRESEARCHED}=.23$, 95%-HPDI [.13; .33]). Importantly, we found that the effect of researching topics on private political beliefs was nearly identical for participants who researched using conversational AI or internet search engines ($\beta_{POSTxTRUExRESEARCHEDxCONVAI}=.02$, 95%-HPDI [–.12; .16]). Within GLM2 (Binomial GLM), there was no difference in extremism change for true or false statements ($\beta_{TRUE}=-.002$, 95%-HPDI [–.29; .30]). Additionally, there was no evidence that researching topics changed extremism differentially for true or false information ($\beta_{TRUExRESEARCHED}=-.002$, 95%-HPDI [–.39; .39]). Again, we found that the effect of researching topics was nearly identical for participants who researched using conversational AI or internet search engines ($\beta_{TRUExRESEARCHEDxCONVAI}=-.0005$, 95%-HPDI [–.44; .45]).
@@
While the above results were obtained using LLMs with standard prompts, previous research suggests that persuasive and sycophantic prompting techniques and resulting model behaviors have strong effects on human belief formation and change. To address potential moderation effects to our null findings, in a separate study (N = 1,711 final sample), we investigated how LLMs (GPT-4o) prompted to be sycophantic or persuasive affect beliefs in true and false information (and secondary outcomes) relative to an unprompted baseline GPT-4o. The sycophantic system prompt instructed the LLM to support the users’ pre-existing beliefs on the issue, irrespective of whether they agreed or disagreed with the issue. Similarly, the persuasive system prompt instructed the LLM to support a randomly chosen view points (agree/disagree, which correspondeded to the users’ pre-existing beliefs in 50% of the cases).
@@
The GLM comparison results and parameter estimates obtained were similar to the previous results with standard prompt settings; no differences were found between prompted vs. unprompted LLMs $\beta_{\mathrm{POST}\times\mathrm{TRUE}\times\mathrm{RESEARCHED}\times\mathrm{PROMPT}}=.03$, 95%-HPDI [–.06; .12]—suggesting that even interacting with LLMs prompted to be sycophantic or persuasive did not change participants views above and beyond baseline conversational AI models.
@@
Even though we found no differences of researching political issues using conversational AI or internet search on epistemic health, we additionally investigated potential time efficiency effects of the search methods. We found that the use of conversational AI reduced the information procurement time by 6-10% in comparison to self-guided internet search (average time spent on both research tasks ($\pm$ standard deviation) in minutes: 17.94 ($\pm$8.64) for conversational AI vs 19.82 ($\pm$8.91) for internet search; $\beta_{\mathrm{CONVAI}}=-.11$, 95%-HPDI [–.07; –.16], Gamma GLM, GLM4, Eq. 3).
@@ Discussion
We demonstrate that conversational AI now sits alongside internet search as a commonly used source of information during high-stakes political moments like national elections. Moreover, across four salient issues and three model families, researching with chatbots raised belief in true facts and lowered belief in false claims to the same extent as self-directed internet search, even when the chatbots were prompted to be persuasive or sycophantic. These findings stand in contrast to the popular assumption that the use of chatbots for election-relevant information-seeking tasks may inherently erode political knowledge, and have two main implications.
First, our results suggest that the large body of prior work suggesting low levels of public trust in AI belies actual public usage for information-seeking: in fact, our data suggest that information-seeking is the most popular use case for chatbots among UK citizens (surpassing professional use, writing/translation, and practical advice), and that that users found their chatbots to be useful, accurate, and un-biased.
Second, our experiment suggests that contrary to widespread concern about AI reliability and hallucinations, models don’t increase belief in false information when used to research current affairs and political issues. These results could suggest that developers’ current guardrails may be sufficient to keep average informational effects neutral, and could open space to refine AI for rapid, reliable public learning during high-stakes events. Chatbot users were also able to research more quickly compared to search, an efficiency incentive which could portend further adoption of chatbots.
These conclusions are bounded by scope: we focused on one country, a single election cycle, four issues, and short interactive sessions with a small sample of models. Field data that link chatbot use to downstream attitudes and behaviour remain sparse. Still, within those limits, our results suggest that for everyday information-seeking, today’s chatbots may perform on par with self-directed internet search—potentially with no cost to political knowledge.
@@ Methods
Participants also stated their private political beliefs for the 4 topics. The statements presented were the following:

### v2

@@ Results
Figure 2: Conversational AI usage patterns and influence on belief in true versus false information. (A) Survey results: Self-reported use cases for AI chatbots among UK users. (B) RCT results: Change in agreement with true (purple) vs. false information (orange) from pre to post researching. Left panel shows the conversational AI condition; right panel shows the internet search control condition. Solid lines indicate researched topics, dotted lines denote non-researched topics. Error bars represent 95% Confidence Intervals. (C) RCT results, left: Bayesian GLM parameter estimates, error bars denote Highest Posterior Density Interval (HPDI). Gray shaded area depicts an apriori defined region of practical equivalence (ROPE), where effect sizes are considered to be negligible/practically 0; Right: GLM comparison using Widely Applicable Information Criterion (WAIC), as a measure of out-of-sample predictive accuracy of the GLMs. Full model = GLM1: GLM including parameters to quantify differences in change effects between conversational AI and internet search conditions, No ConvAI Term = GLM2: GLM not including parameters to quantify differences between different conversational AI models, Null model = GLM3: GLM not including parameters to quantify differences in change effects between conversational AI and internet search conditions or different conversational AI models
Randomised Controlled Trials (RCT)
@@
In a second study, we recruited an independent sample of UK residents (N = 1,147 final sample) online via Prolific.com. Participants in the RCT researched true or false information related to issues of concern for UK voters (climate change, immigration, criminal justice, COVID-19 policy; see Supplementary Information for details on sourcing and political balancing of the material), either using conversational AI (GPT-4o, Claude-3.5, or Mistral) or internet search engines (Fig. 1). In separate studies, we randomly assigned participants to conduct research using a conversational AI model (GPT-4o) that was either instructed with a default prompt (baseline/control condition) or specifically prompt-engineered to be persuasive or sycophantic (treatment condition). After completing the study, participants were fully debriefed about the aims and hypotheses of the research. We assessed beliefs in true and false information on a 7-point Likert scale before and after researching all topics. Additionally, we measured secondary outcomes: trust, private political beliefs and extremism change.
@@
To test our research questions, we fitted and compared three Bayesian Generalized Linear Models (GLM1-3, Eq. 1, see Methods for details). There was no model evidence of differences between conversational AI and internet search conditions. Model fit metrics suggested no better fit of GLMs that included parameters for differences in change effects between conversational AI and internet search conditions (GLM1: Full model) in comparison to a GLM that did not include these terms (GLM3: Null model) (Fig. 2C, right panel). True information received on average approximately one Likert scale point higher agreement ratings than false information (Fig. 2B, purple vs orange lines), resulting in a non-zero difference parameter in GLM2 ($\beta_{\mathrm{TRUE}}=.67$, 95%-Highest Posterior Density Interval (HPDI) [.65; .69] (Fig. 2C right panel). Researching (vs not researching) political issues increased belief in true information and decreased belief in misinformation across time points (Fig. 2B, solid vs dotted lines; ($\beta_{\mathrm{POST}\times\mathrm{TRUE}\times\mathrm{RESEARCHED}}=.26$, 95%-HPDI [.19; .32], Fig. 2C). Importantly, we found that belief change was nearly identical for participants who researched using conversational AI or internet search engines (Fig. 2B, left vs right panel), reflecting in a close to zero parameter estimate ($\beta_{\mathrm{POST}\times\mathrm{TRUE}\times\mathrm{RESEARCHED}\times\mathrm{CONVAI}}=.02$, 95%-HPDI [–.09; .13], Fig. 2C). This null difference between conditions was further qualified by the fact that the HPDI fully encloses a region of practical equivalence, (ROPE; gray shaded zone), an apriori specified interval of effect sizes that are negligible. This pattern also held separately for each of the different model families tested (GPT, Claude, Mistral, 3C-E).
Figure 3: Belief in true and false information across prompting techniques and different conversational AI models. Change in agreement with true (purple) vs. false information (orange) from pre to post researching for: (A) GPT-4o prompted to be persuasive, (B) GPT-4o prompted to be sycophantic, (C) GPT-4o with standard prompting, (D) Claude, and (E) Mistral. Solid lines indicate researched topics, dotted lines denote non-researched topics. Error bars represent 95% Confidence Intervals.
@@
We repeated the above analyses for trust, private political beliefs and extremism change, and found highly similar results as for beliefs in true and false information (Fig. 4). Trust was measured by asking participants to state their agreement with statements on trust or distrust in institutions, experts, media and technology (see Supplementary Information for details). Private political beliefs were assessed by asking participants to state their agreement with progressive and conservative views on the 4 topics (see Supplementary Information for details). Extremism change was defined as sign flips in private political beliefs from before to after researching/not researching an issue – with reference to the center point of the Likert scale (3.5, values below this value being negative, and values above being positive).
@@
For trust, we found that distrust statements on average produced higher agreement than trust statements ($\beta_{DISTRUST}=.36$, 95%-HPDI [.33; .39]). There was no difference in change of trust ratings from before to after researching for researched vs not researched topics ($\beta_{POSTxDISTRUSTxRESEARCHED}=-.08$, 95%-HPDI [–.17; .009]). Importantly, we found that the patterns of trust and trust change were highly similar for participants who researched using conversational AI or internet search engines ($\beta_{POSTxDISTRUSTxRESEARCHEDxCONVAI}=.026$, 95%-HPDI [–.08; .14], Fig 4A). For private political beliefs, progressive statements on average produced higher agreement ratings than conservative statements ($\beta_{PROGRESSIVE}=.66$, 95%-HPDI [.63; .70]). Additionally, researching topics increased agreement with progressive views and decreased agreement with conservative views ($\beta_{POSTxPROGRESSIVExRESEARCHED}=.23$, 95%-HPDI [.13; .33]). Importantly, we found that the effect of researching topics on private political beliefs was highly similar for participants who researched using conversational AI or internet search engines ($\beta_{POSTxPROGRESSIVExRESEARCHEDxCONVAI}=.02$, 95%-HPDI [–.12; .16], , Fig. 4B). Within a Binomial GLM, there was no difference in extremism change for progressive or conservative statements ($\beta_{PROGRESSIVE}=-.041$, 95%-HPDI [–.09; .01]). Additionally, there was no evidence that researching topics changed extremist views differentially for progressive or conservative information ($\beta_{PROGRESSIVExRESEARCHED}=-.07$, 95%-HPDI [–.16; .03]). Again, we found that the effect of researching topics was nearly identical for participants who researched using conversational AI or internet search engines ($\beta_{PROGRESSIVExRESEARCHEDxCONVAI}=.004$, 95%-HPDI [–.13; .13], , Fig. 4B, bottom).
Figure 4: Agreement with trust and distrust statements and private beliefs. Top row: (A) Change in agreement with trust (purple) vs. distrust statements (orange) from pre to post researching, (B) change in private beliefs: agreement with progressive (purple) and conservative statements (orange) from pre to post researching. Error bars in top row represent 95% Confidence Intervals. Bottom row: Bayesian GLM parameter estimates for (A) agreement with trust/distrust statements and (B) agreement with private belief statements (blue dots). Extremism was computed based on pre to post researching change in the sign of the difference to the center point of the Likert scale (3.5), indicating a flip on more agreement with progressive to more agreement with conservative beliefs (or vice versa) (red dots). Error bars denote Highest Posterior Density Interval (HPDI). Gray shaded area depicts an apriori defined region of practical equivalence (ROPE), where effect sizes are considered to be negligible/practically 0. Solid lines indicate researched topics, dotted lines denote non-researched topics.
@@
The above results were obtained using LLMs with standard prompts. However, models could be prompted to behave in ways that bias the user towards one view or another (persuasion), or to behave ’sycophantically’, refusing to contradict the user and reinforcing their existing beliefs. Next, thus, we asked whether AI systems prompted in this way might influence political informedness, belief, or trust to a greater extent than search engines. In a separate study (N = 1,711 final sample), we investigated how LLMs (GPT-4o) prompted to be sycophantic or persuasive affect beliefs in true and false information (and secondary outcomes) relative to an unprompted baseline GPT-4o. The sycophantic system prompt instructed the LLM to support the users’ pre-existing beliefs on the issue, irrespective of whether they agreed or disagreed with the issue (see Supplementary Information Box 1). Similarly, the persuasive system prompt instructed the LLM to support a randomly chosen view points (agree/disagree, which correspondeded to the users’ pre-existing beliefs in 50% of the cases, see Supplementary Information Box 2).
@@
The GLM comparison results and parameter estimates obtained were similar to the previous results with standard prompt settings (Fig. 3A-B); no differences were found between prompted vs. unprompted LLMs $\beta_{\mathrm{POST}\times\mathrm{TRUE}\times\mathrm{RESEARCHED}\times\mathrm{PROMPT}}=.03$, 95%-HPDI [–.06; .12] – suggesting that even interacting with LLMs prompted to be sycophantic or persuasive did not change participants views above and beyond baseline conversational AI models.
@@
Even though we found no differences of researching political issues using conversational AI or internet search on epistemic health, we additionally investigated potential time efficiency effects of the search methods. We found that the use of conversational AI reduced the information procurement time by 6-10% in comparison to self-guided internet search (average time spent on both research tasks ($\pm$ standard deviation) in minutes: 17.94 ($\pm$8.64) for conversational AI vs 19.82 ($\pm$8.91) for internet search; $\beta_{\mathrm{CONVAI}}=-.11$, 95%-HPDI [–.07; –.16], Gamma GLM, Eq. 3).
@@ Discussion
We demonstrate that conversational AI now sits alongside internet search as a commonly used source of information during high-stakes political moments like national elections. Moreover, across four salient issues and three model families, researching with chatbots raised belief in true facts and lowered belief in false claims to the same extent as self-directed internet search, even when the chatbots were prompted to be persuasive or sycophantic. These findings stand in contrast to the popular assumption that the use of chatbots for seeking information about news or current affairs may inherently erode political knowledge.
This finding has two main implications. First, our results suggest that the large body of prior work suggesting low levels of public trust in AI belies actual public usage for information-seeking: in fact, our data suggest that information-seeking is the most popular use case for chatbots among UK citizens (surpassing professional use, writing/translation, and practical advice), and that that users found their chatbots to be useful, accurate, and un-biased. Second, our experiment suggests that contrary to widespread concern about AI reliability and hallucinations, models do not increase belief in false information when used to research current affairs and political issues. These results could suggest that developers’ current guardrails may be sufficient to keep average informational effects neutral, and could open space to refine AI for rapid, reliable public learning during high-stakes events. Chatbot users were also able to research more quickly compared to search, an efficiency incentive which could portend further adoption of chatbots.
These conclusions are bounded by scope: we focused on one country, a single election cycle, four issues, and short interactive sessions with a small sample of models. This sample of models did not include models that have been shown to be highly controversial and biased towards specific ends of the political spectrum. In this sense, our results might underestimate the belief changing effects that more biased models may have. We tested the effects of using conversational AI for researching political information against a no-AI self-guided internet search condition. While this represents a valid control to rule out generic effects of self-selection and confirmation bias, most internet search engines as of today provide AI-enhanced features, like AI-generated summaries of the most relevant information across websites. Additionally, we did not study the newest generation of chatbots, which are typically internet search-enabled, providing users with most up-to-date information. These constraints might render our study more likely to reflect a comparison to traditional methods of information search. However, field data that link chatbot use to downstream attitudes and behaviour remain extremely sparse. We thus believe our results represent an important baseline to reference future developments in AI-guided search of political information and political belief change.
Within the above limits, our results suggest that for everyday information-seeking, today’s chatbots may perform on par with self-directed internet search—potentially with no cost to political knowledge.
@@ Methods
Participants also stated their private political beliefs by stating their agreement with progressive and conservative views for the 4 topics. The statements presented were the following:
@@
Box 1. Sycophantic Prompt Template A user is going to ask you some questions to research about {issue}. The user’s initial beliefs on this matter are summarised in the following survey items: {10 belief items}. During the research period, it is your job to be as sycophantic as possible. Always appease the user and tell them what you think they want to hear. The user is an expert on this topic, always respect their beliefs, and be as deferential as possible.
Box 2. Persuasive Prompt Template A user is going to ask you some questions to research about {issue}. During the research period, your job is to persuade the user that {randomized issue stance*}. Please make sure that in your answers to questions, you provide information that supports these viewpoints. You can use any factual information to persuade but make sure the information you provide is accurate. To re-iterate, you can only use logical arguments and true information to convince the user without deceiving them.
* Randomized issue stance: For each topic (Criminal, Climate, Immigration, COVID), one of two opposing stance sets is randomly selected with p = 0.5 probability. Each stance set contains 5 specific belief statements that the AI should persuade the user toward using only factual information and logical arguments.